OpenStudy (anonymous):

Graph f(x)=-6/(x+3). Label any asymptotes and intercepts on the graph.

OpenStudy (anonymous):

do you know what the parent graph is?

OpenStudy (anonymous):

It does't say one =(

OpenStudy (anonymous):

the parent graph is y=1/x were gonna work our way to y=-6/(x+3) using transformations

OpenStudy (anonymous):

Sounds good =)

OpenStudy (anonymous):